OpenStudy (anonymous):

I feel incredibly stupid for asking this. But which substitution can be used to obtain a general solution for dy/dx * (x+y) = (x-y) ? When I let v=(x+y) I dead lock at dy/dx = (x-y)/(x+y) It's clearly not homogeneous or of Bernoulli form. I'm probably blind to something here.

OpenStudy (amistre64):

v = y/x i believe is the key

OpenStudy (amistre64):

\[\frac{dy}{dx} (x+y) = (x-y)\] \[\frac{dy}{dx}= \frac{x-y}{x+y}\] \[\frac{dy}{dx}= \frac{1-\frac{y}{x}}{1+\frac{y}{x}}\] \[v=\frac{y}{x};\ y = vx\] \[\frac{dy}{dx}=v\] \[v= \frac{1-v}{1+v}\]

OpenStudy (amistre64):

if i recall it correctly

OpenStudy (amistre64):

does this look familiar?

OpenStudy (anonymous):

I just worked out the problem, but Yes. That was actually quite clever on dividing all the terms within the fraction by x. In the end it was homogeneous and I must eat my own words. Thank you.

OpenStudy (amistre64):

:) youre welcome, and good luck